OpenStudy (anonymous):

the question is how do you change the fraction 2/7% to a ratio

OpenStudy (anonymous):

see it is 2: 0.07 But since we do not use decimal in ratios, we can write it as 200 : 7 (Multiplying both by 100 to remove decimal point frm 0.07) Pls check with yr teacher for final verdict

OpenStudy (anonymous):

2/7% = 2/7 x 1/100 = 2/700 = 1/350 so ratio is 1:350

OpenStudy (anonymous):

one to three hundred fifty?

OpenStudy (anonymous):

2 is 200% so harkirat has it i believe. then again i am often wrong

OpenStudy (anonymous):

I think the question is 2 : 7% because it says that it is a fraction

OpenStudy (anonymous):

yeah then its ambiguous..i thought (2/7)%

OpenStudy (anonymous):

or : 0.07= 7/100 2/7/100=2*(100/7)

OpenStudy (anonymous):

Yeah that is another way of looking at it Nancy...

OpenStudy (anonymous):

what is the question guys? lets get it clear from ibtburns

OpenStudy (anonymous):

the question is change the fraction 2/7% into a ratio

OpenStudy (anonymous):

hey bhagwan! is it (2/7)% or 2 / (7%)????

OpenStudy (anonymous):

\[\frac{2}{0.07}=28.57142857\]or\[0.07=\frac{7}{100}\]\[\frac{2}{\frac{7}{100}}=2\times\frac{100}{7}=28.57142857\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

u r we come

OpenStudy (anonymous):

Nancy 28.57.... is not a ratio We have to giv ans as a ratio

OpenStudy (anonymous):

no I id not say ratio, I just write the evalue